# An Exploration of Default Images in Text-to-Image Generation

### v5

@@ 3.4. Experiments and User Studies on TTI Quality and Satisfaction
Evaluations of TTI models often rely on automatic metrics and curated benchmarks, yet these measures are known to misalign with human perception (2203.06026; otani2023verifiablereproduciblehumanevaluation). The Fréchet Inception Distance (FID), for instance, can yield scores that disagree with user judgments of image quality and prompt alignment, limiting its utility as a sole evaluation criterion (2203.06026; 2403.11821; otani2023verifiablereproduciblehumanevaluation). Complementary metrics – such as CLIPScore (CLIPScore) – have been proposed to capture semantic fidelity, but also exhibit gaps when compared to subjective assessments (VQAScore). Recognizing these shortcomings, russo-2022-creative suggested criteria for a creative TTI benchmark incorporating insights from linguistics, cognitive psychology, and art theory to better capture human notions of creativity and relevance (russo-2022-creative). More recently, the Holistic Evaluation of Text-to-Image Models (HEIM) benchmark systematically assessed 26 state-of-the-art TTI generators across twelve dimensions (ranging from aesthetics and originality to bias and robustness) using a combination of automated measures and human ratings, providing a richer, multidimensional view of model performance (2311.04287).
@@
Human‐centered studies bridge the gap between automatic scores and actual user experience. The Pick-a-Pic dataset includes over half a million human preference annotations, enabling the training of PickScore, a model predicting human rankings of generated images with superhuman accuracy, and suggesting a more reliable protocol for evaluating TTI outputs (NEURIPS202373aacd8b). Researchers extended this work with ImageReward, a preference‐based reward model from 137,000 expert comparisons. Fine-tuning diffusion models against this reward function produces images that users consistently rate as higher quality (xu2023imagerewardlearningevaluatinghuman). otani2023verifiablereproduciblehumanevaluation highlight the lack of verifiable, reproducible human evaluation in TTI research and propose a standardized, crowdsourced protocol (otani2023verifiablereproduciblehumanevaluation). Their results reveal that automatic metrics frequently diverge from human judgments, underscoring the need for rigorous user studies. In interactive settings, tools, such as PromptAid (mishra2025promptaidpromptexplorationperturbation) and PromptCharm (3613904.3642803), helpedparticipants craft prompts yielding images better aligned with their intent, with users reporting increased satisfaction and perceived control over output quality. Multi-turn guidance systems, such as DialPrompt (2408.12910), achieve significant gains in both image quality ratings and user‐centricity scores compared to baseline prompt‐engineering interfaces. Finally, exploratory work in teacher education demonstrates how practitioners reflect on prompt strategies and generator behaviors in art and design, revealing both learning affordances and cognitive challenges when integrating TTI tools into creative pedagogy (Ringvold).
@@ 4.1. Input Prompts Likely Triggering Default Images
    Glitch tokens (SolidGoldMagikarp) (e.g. rawdownload)
@@
We manually created a set of 130 prompts based on the six categories, A1–A6. We created prompts in category A2 by introducing altering known words with slight variations. We selected the two low-resource languages (A4.1 Finnish and A4.2 Tagalog) for their language-native words that are not borrowed from other languages, so they do not cognate with widely known foreign languages. We collected the glitch tokens from related literature (SolidGoldMagikarp; SolidGoldMagikarp2; SolidGoldMagikarp3). The prompts are in Appendix A).
To assess the likelihood that the selected prompts had been encountered during model training, we checked each prompt on haveibeentrained.com, a search interface for over 5.8 billion image-caption pairs. The tool helps users determine whether a specific term (e.g., their name) appears in large-scale datasets, such as LAION (Laion). These datasets are commonly used in the TTI model training (Laion; NEURIPS2022a1859deb). We sought prompts that were absent from – or underrepresented in – the training data, increasing the likelihood that the resulting generated imagery would involve default or fallback behavior. Our inclusion criteria were twofold: 1) the prompt should plausibly occur in real-world use, and 2) it should pose a challenge to the model due to low training data coverage. We describe our image generation process next.
@@ 4.2. Image Generation Experiments and Ablation Studies
We used Midjourney’s webpage333www.midjourney.com/imagine (see Figure 2) to execute the experiments, using fixed parameters listed in Table 1. For each of the 130 prompts in the six input categories (A1–A6), we ran 1 trial, resulting in four images per prompt, for a total of 520 images. We then visually analyzed the images, as described in Section 1.
@@
As a sensitivity analysis to test the robustness of default images, we conducted five small-scale ablation studies. We investigated the effect of different seed values on the default image (Section 5.2.1). We explored whether default images could be altered with prompt modifiers (taxonomy) (Section 5.2.2). We then examined default image behavior within larger prompts, including known terms (Section 5.2.3) and combinations of prompts known to trigger default images (Section 5.2.4. In our last study, we investigated default images in different model versions (Section 5.2.5).
@@ 4.3. Default Image Identification
To identify default images in the set of generated images, we used affinity diagramming (or KJ Method) (KJMethod). It is a qualitative method that organizes data into groups based on perceived similarity. The method is commonly employed to categorize complex data. The first two authors collaboratively grouped images on a digital canvas based on their perceived similarity. During the session, the two authors discussed their rationale and iteratively refined the groups. Simultaneously, the two authors developed descriptive labels for the images (e.g., “Lady-Birdhead”), enabling them to refer to these images by name. They continued this collaborative process until all images were either categorized as default images (those with strong similarities across multiple images created with different prompts) or not. Identifying similarities between images was often straightforward despite subtle variations, as visible in Figure 5.
@@ 4.4. Online User Study
We used crowdsourcing through Prolific,444www.prolific.com an online human subject pool, to explore participant perception of default images, specifically, the impact of default images on user satisfaction. We used an onlinea questionnaire presented participants with four images, including the original prompt that created the image. Unknown to the participant, for two of the images (Q1 and Q2), we replaced one term in the prompt with a term identified as causing default images. For instance, in Q1, we replaced ‘raven’ in the prompt “photoshoot of a woman wearing a white off-shoulder dress with a black raven on top of her head” with ’Greagoft’. For Q2, we replaced “leaves and flowers” with “Shuttastfr”. With this replacement, we simulated the situation where a word is unknown to the generative model. We showed participants the original prompt and the image resulting from the modified prompt, simulating the mismatch in expectations that an encounter with a default image can produce. For Q3, we used an image that matched the descriptive prompt, without replacements. For Q4, we used a prompt that was unknown to the model and, thus, the accompanying image was a default image. The four prompts and images are in Figure 8. We asked participants to rate their satisfaction with each image, given the prompt, on a scale from 1 (Strongly Dissatisfied) to 7 (Strongly Satisfied). Participants received £6.29 per hour, and the median time to complete the survey was 3 min and 49 seconds.
@@ 4.5.2. Dataset description
The dataset contains 189,431 prompts, from July 9, 2022 to March 17, 2025. Each prompt produced exactly four images, for a total of 757,724 images. This includes 29,868 images for Midjourney version 1.0, 92,496 images for v2.0, 173,468 images for v3.0, 109,672 images for v4.0, 64,092 images for v5.0, 67,248 images for v5.1, 60,300 images for v5.2, 108,868 images for v6.0, and 51,712 images for v6.1.The mean prompt length is 17.5 characters ($Median=18$, $SD=5.5$, $Q1=14$, $Q3=22$). Some prompts include special tokens and punctuation, although single character prompts are rare (see Figure 3). The prompt languages include English (47.2%), Italian (5.4%), Romanian (4.0%), French (3.1%), and many others. The columns included in our dataset are described in Table 2 in Appendix B.
@@ 4.5.3. Data analysis
We experimented with different solutions to identify visually similar images, including training a joint image-prompt embedding with a contrastive learning loss, as well as image similarity measures, such as cosine similarity between pixel vectors, LPIPS (1801.03924.pdf), SSIM (1284395), and MS-SSIM (1284396). We found the CLIP model (ViT-L/14) (CLIP; VIT) produced the most visually coherent clusters for our use case. Using CLIP, we transformed the images into embeddings (i.e., numeric vectors), which we precomputed locally and then transferred to a cloud server. We then computed the full pairwise cosine similarity matrix between embeddings to identify visually similar images. Because the number of unique pairs grows quadratically with the number of images, the computation and storage scale as $O(N^{2})$. In practice, we exploit symmetry by evaluating only the upper triangle and mirroring. To fit in memory, we computed matrices separately for each major Midjourney version. All computations were performed on a cloud server with 96 vCPUs and 1 TB RAM.
To find clusters among the visually similar images, we employed a clustering algorithm. Again, we experimented with different solutions, such as connected component clustering and HDBSCAN (HDBSCAN), but settled on simple agglomerative hierarchical clustering (10.1002/widm.53) with a threshold of 0.1 on cosine distances between normalized CLIP embeddings. This threshold controls cluster formation by stopping merges once the cosine distance between clusters exceeds 0.1, thereby grouping images with high visual similarity. We set the minimum cluster size to $n=3$ images. From visual inspection, we noted that clusters were individually tight but sometimes similar to each other across clusters, a common artifact of hierarchical clustering with a fixed distance threshold. To address this, we implemented a post-processing merge of clusters based on their centroid similarity with a threshold of 0.9.
@@ 5.1. Default Images
The default images we identified exhibit recurring motifs, reflecting the model’s aesthetic biases when faced with ambiguous or unrecognized prompts. Human figures and natural elements – such as birds, clouds, trees, and flowers – appear as primary motifs, separately or merged into strange creatures. Human-vegetal and/or human-animal fusions emerge as key subjects in default images, appearing across different atmospheres, from oneiric scenes to mythological metamorphoses. Some identified default images confirm that similarly to human dreams, “AI has an uncanny ability to combine seemingly disparate elements into sometimes sinisterly coherent representations” (PRADA).
@@
Although the literature on AI imagery attests that visual expression tends to vary according to specific TTI generators, and some note ‘default styles’ associated with different models (MANOVICH), the curated series of 10 images presented in Figure 4 reveals significant aesthetic disparities. While the five images in the lower row are reminiscent of abstract expressionist and surrealist painting, also evoking magical realism and epic fantasy films, the first five images draw from other artistic movements. As shown in figures 4 and 5, the female portrait is also an important theme, particularly prone to formal and stylistic derivations. From the solitary woman standing before a romantic landscape – in which Caspar David Friedrich’s painting resonates – to close-up representations that seem to combine the tradition of portrait painting with lyrical abstraction and contemporary fashion photography, these default images illustrate the hybrid and composite nature of AI-generated pictures. Figure 5 further emphasizes how AI-generated pictures can condense multiple artistic references into a single visual composition that, similarly to a musical piece, can then unfold in subtle variations. This series reflects the wide dissemination of individual portraits, both through the digitization and online circulation of art collections – including popular museum masterpieces – and through the massive production of born-digital pictures that fuel social media.
@@
The intriguing association of a stereotyped female figure and a bird in place of a hat in Figure 5 poses a twofold question concerning the generation of default images. Apparently, the TTI model was unable to understand the prompt or correctly recognize some of the images it was trained with. Feather hats are a recurring theme in the history of female portraiture and fashion since the Renaissance, from Titian’s Portrait of a Young Woman (c. 1536) to Rubens’ Portrait of Helena Fourment (c. 1630), from Klimt’s The Black Feathered Hat (1910) to Karl Lagerfeld’s exuberant designs for Chanel’s Spring Collection in 1992. The examples are countless, hence it seems plausible that many of these images were used to train the model. However, there seems to be a fundamental difference between human perception of visual representations of feather hats and the way computer vision ‘sees’ such images. Rather than a dataset bias, what is at stake here is a perceptual bias (OFFERT). Following a metonymical logic (i.e., taking a part to stand in for the whole), the model takes the part for the whole, assuming that the feathers correspond to the presence of a bird whose species differs across the variations of the default image. This is an insightful demonstration of “the nomadic existence to which images seem to be condemned within the historical formations of knowledge” (ALLOA).
5.2. Ablation Studies
5.2.1. Effects of seed values
@@ 6.1. Default Images in Text-to-Image Generation
Our computational analysis of images on Midjourney reveals that default images are not a theoretical construct, they occur daily in response to real-world creative user prompts. Some of the images we identified in our analysis represent iconic Midjourney failure modes that many users may have encountered, especially with the early Midjourney versions. Our findings confirm that these default images are not generated randomly and, on the basis of our systematic experiments, have a relatively narrow thematic and formal range. Uncanny elements emerged in common motifs that often depicted human portraits (usually women) and animals (especially birds), along with other botanical or organic forms. Earlier versions of Midjourney have a different set of default images, often including motifs such as balloons, butterflies, fishes, floating islands, towns, or simply color pallettes and transitions (see Figure 1 and Appendix C). Although many default images appear to share a common style, this straightforward association may obfuscate a more complex reality. For instance, in model version 5, some default images appear related to surrealism and cartoonish grotesque (cf. Figure 1), with other identifiable references from art history, namely Dutch and Flemish Renaissance painting, particularly Hieronymus Bosch and Bruegel the Elder. AI-generated images are hybrid and composite, condensing multiple forms and references learned from the training dataset into a single visual output. As Wasielewski points out “style is a gradient” (Wasielewski). Therefore, an AI image is unlikely to correspond to a specific artistic movement, rather reflecting “a fluctuating set of visual characteristics with highly fuzzy boundaries” (Wasielewski).

### v6

@@ 3.4. Experiments and User Studies on TTI Quality and Satisfaction
Evaluations of TTI models often rely on automatic metrics and curated benchmarks, yet these measures are known to misalign with human perception (Kynkäänniemi et al., 2023; Otani et al., 2023). The Fréchet Inception Distance (FID), for instance, can yield scores that disagree with user judgments of image quality and prompt alignment, limiting its utility as a sole evaluation criterion (Kynkäänniemi et al., 2023; Hartwig et al., 2025; Otani et al., 2023). Complementary metrics – such as CLIPScore (Hessel et al., 2021) – have been proposed to capture semantic fidelity, but also exhibit gaps when compared to subjective assessments (Lin et al., 2024). Recognizing these shortcomings, Russo suggested criteria for a creative TTI benchmark incorporating insights from linguistics, cognitive psychology, and art theory to better capture human notions of creativity and relevance (Russo, 2022). More recently, the Holistic Evaluation of Text-to-Image Models (HEIM) benchmark systematically assessed 26 state-of-the-art TTI generators across twelve dimensions (ranging from aesthetics and originality to bias and robustness) using a combination of automated measures and human ratings, providing a richer, multidimensional view of model performance (Lee et al., 2023).
@@
Human‐centered studies bridge the gap between automatic scores and actual user experience. The Pick-a-Pic dataset includes over half a million human preference annotations, enabling the training of PickScore, a model predicting human rankings of generated images with superhuman accuracy, and suggesting a more reliable protocol for evaluating TTI outputs (Kirstain et al., 2023). Researchers extended this work with ImageReward, a preference‐based reward model from 137,000 expert comparisons. Fine-tuning diffusion models against this reward function produces images that users consistently rate as higher quality (Xu et al., 2023). Otani et al. highlight the lack of verifiable, reproducible human evaluation in TTI research and propose a standardized, crowdsourced protocol (Otani et al., 2023). Their results reveal that automatic metrics frequently diverge from human judgments, underscoring the need for rigorous user studies. In interactive settings, tools, such as PromptAid (Mishra et al., 2025) and PromptCharm (Wang et al., 2024), helpedparticipants craft prompts yielding images better aligned with their intent, with users reporting increased satisfaction and perceived control over output quality. Multi-turn guidance systems, such as DialPrompt (Liu et al., 2024), achieve significant gains in both image quality ratings and user‐centricity scores compared to baseline prompt‐engineering interfaces. Finally, exploratory work in teacher education demonstrates how practitioners reflect on prompt strategies and generator behaviors in art and design, revealing both learning affordances and cognitive challenges when integrating TTI tools into creative pedagogy (Ringvold et al., 2023).
@@ 4.1. Input Prompts Likely Triggering Default Images
    Glitch tokens (Rumbelow and Watkins, 2023) (e.g. rawdownload)
@@
We manually created a set of 130 prompts based on the six categories, A1–A6. We created prompts in category A2 by introducing altering known words with slight variations. We selected the two low-resource languages (A4.1 Finnish and A4.2 Tagalog) for their language-native words that are not borrowed from other languages, so they do not cognate with widely known foreign languages. We collected the glitch tokens from related literature (Rumbelow and Watkins, 2023; Watkins and Rumbelow, 2023a, b). The prompts are in Appendix A).
To assess the likelihood that the selected prompts had been encountered during model training, we checked each prompt on haveibeentrained.com, a search interface for over 5.8 billion image-caption pairs. The tool helps users determine whether a specific term (e.g., their name) appears in large-scale datasets, such as LAION (LAION.ai, n.d.). These datasets are commonly used in the TTI model training (LAION.ai, n.d.; Schuhmann et al., 2022). We sought prompts that were absent from – or underrepresented in – the training data, increasing the likelihood that the resulting generated imagery would involve default or fallback behavior. Our inclusion criteria were twofold: 1) the prompt should plausibly occur in real-world use, and 2) it should pose a challenge to the model due to low training data coverage. We describe our image generation process next.
@@ 4.2. Image Generation Experiments and Ablation Studies
We used Midjourney’s webpage222www.midjourney.com/imagine (see Figure 2) to execute the experiments, using fixed parameters listed in Table 1. For each of the 130 prompts in the six input categories (A1–A6), we ran 1 trial, resulting in four images per prompt, for a total of 520 images. We then visually analyzed the images, as described in Section 1.
@@
As a sensitivity analysis to test the robustness of default images, we conducted five small-scale ablation studies. We investigated the effect of different seed values on the default image (Section 5.2.1). We explored whether default images could be altered with prompt modifiers (Oppenlaender, 2023) (Section 5.2.2). We then examined default image behavior within larger prompts, including known terms (Section 5.2.3) and combinations of prompts known to trigger default images (Section 5.2.4. In our last study, we investigated default images in different model versions (Section 5.2.5).
@@ 4.3. Default Image Identification
To identify default images in the set of generated images, we used affinity diagramming (or KJ Method) (Scupin, 1997). It is a qualitative method that organizes data into groups based on perceived similarity. The method is commonly employed to categorize complex data. The first two authors collaboratively grouped images on a digital canvas based on their perceived similarity. During the session, the two authors discussed their rationale and iteratively refined the groups. Simultaneously, the two authors developed descriptive labels for the images (e.g., “Lady-Birdhead”), enabling them to refer to these images by name. They continued this collaborative process until all images were either categorized as default images (those with strong similarities across multiple images created with different prompts) or not. Identifying similarities between images was often straightforward despite subtle variations, as visible in Figure 5.
@@ 4.4. Online User Study
We used crowdsourcing through Prolific,333www.prolific.com an online human subject pool, to explore participant perception of default images, specifically, the impact of default images on user satisfaction. We used an onlinea questionnaire presented participants with four images, including the original prompt that created the image. Unknown to the participant, for two of the images (Q1 and Q2), we replaced one term in the prompt with a term identified as causing default images. For instance, in Q1, we replaced ‘raven’ in the prompt “photoshoot of a woman wearing a white off-shoulder dress with a black raven on top of her head” with ’Greagoft’. For Q2, we replaced “leaves and flowers” with “Shuttastfr”. With this replacement, we simulated the situation where a word is unknown to the generative model. We showed participants the original prompt and the image resulting from the modified prompt, simulating the mismatch in expectations that an encounter with a default image can produce. For Q3, we used an image that matched the descriptive prompt, without replacements. For Q4, we used a prompt that was unknown to the model and, thus, the accompanying image was a default image. The four prompts and images are in Figure 8. We asked participants to rate their satisfaction with each image, given the prompt, on a scale from 1 (Strongly Dissatisfied) to 7 (Strongly Satisfied). Participants received £6.29 per hour, and the median time to complete the survey was 3 min and 49 seconds.
@@ 4.5.2. Dataset description
The dataset contains 189,432 prompts, from July 9, 2022 to March 17, 2025. Each prompt produced exactly four images, for a total of 757,728 images. This includes 29,872 images for Midjourney version 1.0, 92,496 images for v2.0, 173,468 images for v3.0, 109,672 images for v4.0, 64,092 images for v5.0, 67,248 images for v5.1, 60,300 images for v5.2, 108,868 images for v6.0, and 51,712 images for v6.1.The mean prompt length is 17.5 characters ($Median=18$, $SD=5.5$, $Q1=14$, $Q3=22$). Some prompts include special tokens and punctuation, although single character prompts are rare (see Figure 3). The prompt languages include English (47.2%), Italian (5.4%), Romanian (4.0%), French (3.1%), and many others. The columns included in our dataset are described in Table 2 in Appendix B.
@@ 4.5.3. Data analysis
We experimented with different solutions to identify visually similar images, including training a joint image-prompt embedding with a contrastive learning loss, as well as image similarity measures, such as cosine similarity between pixel vectors, LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and MS-SSIM (Jalobeanu et al., 2004). We found the CLIP model (ViT-L/14) (Radford et al., 2021; Dosovitskiy et al., 2021) produced the most visually coherent clusters for our use case. Using CLIP, we transformed the images into embeddings (i.e., numeric vectors), which we precomputed locally and then transferred to a cloud server. We then computed the full pairwise cosine similarity matrix between embeddings to identify visually similar images. Because the number of unique pairs grows quadratically with the number of images, the computation and storage scale as $O(N^{2})$. In practice, we exploit symmetry by evaluating only the upper triangle and mirroring. To fit in memory, we computed matrices separately for each major Midjourney version. All computations were performed on a cloud server with 96 vCPUs and 1 TB RAM.
To find clusters among the visually similar images, we employed a clustering algorithm. Again, we experimented with different solutions, such as connected component clustering and HDBSCAN (Campello et al., 2013), but settled on simple agglomerative hierarchical clustering (Murtagh and Contreras, 2012) with a threshold of 0.1 on cosine distances between normalized CLIP embeddings. This threshold controls cluster formation by stopping merges once the cosine distance between clusters exceeds 0.1, thereby grouping images with high visual similarity. We set the minimum cluster size to $n=3$ images. From visual inspection, we noted that clusters were individually tight but sometimes similar to each other across clusters, a common artifact of hierarchical clustering with a fixed distance threshold. To address this, we implemented a post-processing merge of clusters based on their centroid similarity with a threshold of 0.9.
@@ 5.1. Default Images
The default images we identified exhibit recurring motifs, reflecting the model’s aesthetic biases when faced with ambiguous or unrecognized prompts. Human figures and natural elements – such as birds, clouds, trees, and flowers – appear as primary motifs, separately or merged into strange creatures. Human-vegetal and/or human-animal fusions emerge as key subjects in default images, appearing across different atmospheres, from oneiric scenes to mythological metamorphoses. Some identified default images confirm that similarly to human dreams, “AI has an uncanny ability to combine seemingly disparate elements into sometimes sinisterly coherent representations” (Martín Prada, 2025).
@@
Although the literature on AI imagery attests that visual expression tends to vary according to specific TTI generators, and some note ‘default styles’ associated with different models (Manovich, 2024), the curated series of 10 images presented in Figure 4 reveals significant aesthetic disparities. While the five images in the lower row are reminiscent of abstract expressionist and surrealist painting, also evoking magical realism and epic fantasy films, the first five images draw from other artistic movements. As shown in figures 4 and 5, the female portrait is also an important theme, particularly prone to formal and stylistic derivations. From the solitary woman standing before a romantic landscape – in which Caspar David Friedrich’s painting resonates – to close-up representations that seem to combine the tradition of portrait painting with lyrical abstraction and contemporary fashion photography, these default images illustrate the hybrid and composite nature of AI-generated pictures. Figure 5 further emphasizes how AI-generated pictures can condense multiple artistic references into a single visual composition that, similarly to a musical piece, can then unfold in subtle variations. This series reflects the wide dissemination of individual portraits, both through the digitization and online circulation of art collections – including popular museum masterpieces – and through the massive production of born-digital pictures that fuel social media.
@@
The intriguing association of a stereotyped female figure and a bird in place of a hat in Figure 5 poses a twofold question concerning the generation of default images. Apparently, the TTI model was unable to understand the prompt or correctly recognize some of the images it was trained with. Feather hats are a recurring theme in the history of female portraiture and fashion since the Renaissance, from Titian’s Portrait of a Young Woman (c. 1536) to Rubens’ Portrait of Helena Fourment (c. 1630), from Klimt’s The Black Feathered Hat (1910) to Karl Lagerfeld’s exuberant designs for Chanel’s Spring Collection in 1992. The examples are countless, hence it seems plausible that many of these images were used to train the model. However, there seems to be a fundamental difference between human perception of visual representations of feather hats and the way computer vision ‘sees’ such images. Rather than a dataset bias, what is at stake here is a perceptual bias (Offert and Bell, 2021). Following a metonymical logic (i.e., taking a part to stand in for the whole), the model takes the part for the whole, assuming that the feathers correspond to the presence of a bird whose species differs across the variations of the default image. This is an insightful demonstration of “the nomadic existence to which images seem to be condemned within the historical formations of knowledge” (Alloa, 2021).
@@ 6.1. Default Images in Text-to-Image Generation
Our computational analysis of images on Midjourney reveals that default images are not a theoretical construct, they occur daily in response to real-world creative user prompts. Some of the images we identified in our analysis represent iconic Midjourney failure modes that many users may have encountered, especially with the early Midjourney versions. Our findings confirm that these default images are not generated randomly and, on the basis of our systematic experiments, have a relatively narrow thematic and formal range. Uncanny elements emerged in common motifs that often depicted human portraits (usually women) and animals (especially birds), along with other botanical or organic forms. Earlier versions of Midjourney have a different set of default images, often including motifs such as balloons, butterflies, fishes, floating islands, towns, or simply color pallettes and transitions (see Figure 1 and Appendix C). Although many default images appear to share a common style, this straightforward association may obfuscate a more complex reality. For instance, in model version 5, some default images appear related to surrealism and cartoonish grotesque (cf. Figure 1), with other identifiable references from art history, namely Dutch and Flemish Renaissance painting, particularly Hieronymus Bosch and Bruegel the Elder. AI-generated images are hybrid and composite, condensing multiple forms and references learned from the training dataset into a single visual output. As Wasielewski points out “style is a gradient” (Wasielewski, 2023). Therefore, an AI image is unlikely to correspond to a specific artistic movement, rather reflecting “a fluctuating set of visual characteristics with highly fuzzy boundaries” (Wasielewski, 2023).